If the market for bread is described by the following equations Qd=210-0,6P and Qs=-96+1,1P what is the equilibrium price for bread?  

1
Expert's answer
2021-06-02T09:38:12-0400

If the market for bread is described by the following equations: Qd = 210 - 0.6P and Qs = -96 + 1.1P, then the equilibrium price for bread is:

Qd = Qs,

210 - 0.6P = -96 + 1.1P,

1.7P = 306,

P = 180.